# FIVB ranking: Misstep in the right direction

### v1

@@ 5 Real-time ranking
  • The adaptation step $\mu=0.01$ used by the FIVB ranking seems too small and, by increasing it three- or four-fold, performance improves. This can be explained using the interpretation of the SG algorithm as a simplified Kalman filter (Szczecinski and Tihon, 2023, Sec. 3.3), where the adaptation step in the SG algorithm has a meaning of the posterior variance of the skills. In simple terms, the FIVB is over-optimistic about the uncertainty (variance) in the estimation of the skills.
  • Since the weighting may also be interpreted as a variable step size, by removing it, i.e., using $\xi_{v}\equiv 1$, we have to explicitly increase the step size; this explains the large value of $\mu$ for each configuration in which we use $\xi_{v}=1$.
@@ 6 Conclusions
In this work, the online ranking algorithm used by the FIVB is presented in the statistical learning framework. To our best knowledge, the FIVB ranking is the first to adopt an explicit probabilistic model (here, the Cumulative Link (CL)model) of the multi-level ordinal outcomes, and, from the statistical perspective, this is a step in the right direction. On the other hand, the algorithms adopt simplifications which we demonstrate to be suboptimal, but we show how these may be easily corrected.
To analyze the algorithm, we use two approaches: i) the analytical, where the approximations and simplifications allow us to draw conclusions about the properties of the model, as well as to optimize its parameters, and ii) the numerical, where we explicitly optimize the parameters of the model from the outcomes of the international volleyball games used in the FIVB ranking.
The analytical approach is easily reproducible, while the numerical optimization which relies on the cross-validation strategy allows us to validate the insights obtained analytically.
The understanding of the FIVB ranking algorithm may be summarized as follows:
  • The FIVB algorithm should be seen as the approximate ML inference of the skills from the ordinal game outcomes. The approximations are due to the use of the SG to solve the optimization problem, and, more importantly, due to the use of the loss functions, which are proxies for the log-likelihood of the ML approach. We explain the rationale for using such proxy loss functions.
  • Although the form of loss functions is not explicitly mentioned in the description of the FIVB algorithm, they can be inferred from the equations, and we show how they depend on the numerical scores that are attributed to the ordinal game outcomes in the FIVB algorithm. This is interesting because in this way we explain the meaning of the numerical scores, as, from the modeling perspective, the ordinal variables do not have numerical values.
Regarding the model underlying the current FIVB algorithm, we studied its important elements:
  • CL model thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$, see (23), which define the probabilistic ordinal model of the data. They fit relatively well the data we analyzed (FIVB games from 2021-2023), see Sec. 3.2.
@@
  • Importance weights $\boldsymbol{\xi}$, see Table 1, which change the contribution of the outcome of the game depending on the game type. These are shown to be irrelevant from an statistical point of view, see Sec. 3.4.
  • Home-field advantage (HFA) which deals with the games played on the home venues by artificially boosting the skills of the home-team. Here, contrary to the statements in the FIVB descriptions, we found that the HFA is a relevant parameter that improves the prediction performance.
  • CDF function, which defines the CL model, see (21). Replacing the Gaussian CDF used currently by the FIVB ranking with the logistic CDF, see (72), produces multiple benefits: i) the resulting algorithm is very simple, see (82) and (81), ii) it makes the concept of the numerical score superfluous, and iii) its thresholds $\boldsymbol{c}^{\mathcal{L}}$ can be obtained though a simple transformation of the thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$ currently used in the FIVB ranking, see (78).
@@
Since the FIVB explicitly says that its algorithm may be updated, if this is to happen, our recommendations are the following:
@@
    Introduce the HFA to the algorithm. This is a simple and low-cost modification, yielding an improvement in the prediction of the games played on the home-venue.
@@
    Remove the weighting of the games with $\xi^{{\textnormal{FIVB}}}_{v}$. Or, if its use is motivated by some extra-statistical reasons, decrease the differences between the possible values of $\xi_{v}^{{\textnormal{FIVB}}}$.
  4. 4.
    Use the logistic CDF in the CL model, as this removes the need for auxiliary numerical scores (thus making the recommendation #2 void) and yields a simple, and easily interpretable algorithm.
@@ Appendix A Notation in FIVB ranking
We show in Table 6, the relationship between our notation and the one used in the FIVB ranking description FIVB (2024), where the following abbreviations are used
@@ Appendix B Optimization of the cross-validation metric
However, the expressions for the gradient and, especially, the Hessian, quickly become cumbersome; see Burn (2020). Thus, instead of explicit differentiation, we use the automatic differentiation available in JAX and JAXopt python-compliant packages Bradbury et al. (2018), Blondel et al. (2021) with a particularly interesting feature of implicit differentiation required to find the derivative of $\hat{\boldsymbol{\theta}}$ with respect to hyperparameters in $\boldsymbol{p}$ as specified by (105).

### v2

@@ F POL USA JPN GER ARG SLO ITA
  • The adaptation step $\mu=0.01$ used by the FIVB ranking is too small and, by increasing it three- or four-fold, performance improves. This can be explained using the interpretation of the SG algorithm as a simplified Kalman filter proposed by Szczecinski and Tihon (2023, Sec. 3.3), where the adaptation step in the SG algorithm has a meaning of the posterior variance of the skills. In simple terms, FIVB is over-optimistic about the uncertainty (variance) in the estimation of the skills.
  • Since the weighting may also be interpreted as a variable step size, by removing it, i.e., using $\xi_{v}\equiv 1$, we have to explicitly increase the step size; this explains the large value of $\mu$ for each configuration in which we use $\xi_{v}\equiv 1$.
@@ 5 Conclusions
In this work, the online ranking algorithm used by the FIVB is presented in the statistical learning framework. To our best knowledge, the FIVB ranking is the first to adopt an explicit probabilistic model (here, the Cumulative Link (CL)model) of the multi-level ordinal outcomes, and, from the statistical perspective, this is a step in the right direction. On the other hand, the algorithms adopt simplifications that we demonstrate to be suboptimal, which is the “misstep” in the title. However, we show how these simplifications may be easily corrected using well-defined formulas to calculate the numerical scores, see (46). The impact of these changes on the on-line ranking can be seen in Table 4 and Table 5.
To analyze the algorithm, we use two approaches: i) the analytical, where the approximations and simplifications allow us to draw conclusions about the properties of the model, as well as to optimize its parameters, and ii) the numerical, where we explicitly optimize the parameters of the model from the outcomes of the international volleyball matches used in the FIVB ranking.
The analytical approach is easily reproducible, while the numerical optimization which relies on the cross-validation strategy allows us to validate the insights obtained analytically. This led to the following understanding of the current FIVB ranking algorithm:
  • The FIVB algorithm should be seen as the approximate ML inference of the skills from the ordinal match outcomes. The approximations are due to the use of the SG to solve the optimization problem, and, more importantly, due to the use of the loss functions, which are proxies for the log-likelihood of the ML approach. We explain the rationale for using such proxy loss functions.
  • Although the form of loss functions is not explicitly mentioned in the description of the FIVB algorithm, they can be inferred from the equations, and we show how they depend on the numerical scores that are attributed to the ordinal match outcomes in the FIVB algorithm. This is interesting because in this way we explain the meaning of the numerical scores, as, from the modeling perspective, the ordinal variables do not have numerical values.
Regarding the model underlying the current FIVB ranking algorithm and the algorithm itself, we studied:
  • CL model thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$, see (18), which define the probabilistic ordinal model of the data. They fit relatively well the data we analyzed (the FIVB matches from 2021-2023), see Sec. 3.2.
@@
  • Importance weights $\boldsymbol{\xi}$, see Table 1, which change the contribution of the outcome of the match depending on the match type. These are shown to be irrelevant from an statistical point of view, see Sec. 3.4.
  • Home-field advantage (HFA) which deals with the matches played on the home venues by artificially boosting the skills of the home-team. We found that the HFA is a relevant parameter that improves the prediction performance. The improvements are relatively small, which may explain why the current FIVB ranking does not use the HFA. On the other hand, there is no cost related to its application.
@@
In summary, by keeping the structure of the current FIVB ranking and by exploring the optimality of the above model/algorithm choices, we came up with new parameters that can improve the performance of the algorithm. And, since FIVB explicitly says that its algorithm may be updated, if this is to happen, our recommendations, in order of importance, are the following:
@@
    Introduce the HFA to the algorithm. This is a simple and low-cost modification, yielding an improvement in the prediction of the matches played on the home-venue.
@@
    Remove the weighting of the matches with $\xi^{{\textnormal{FIVB}}}_{v}$. Or, if its use is motivated by some extra-statistical (e.g., entertainment) reasons, decrease the differences between the possible values of $\xi_{v}^{{\textnormal{FIVB}}}$.
Further work
While in this work, we focus on the FIVB ranking, the evaluation methodology we propose can be used more broadly, to analyze the ranking algorithms. In particular, the “reverse engineering” approach we used to reveal the form of the (implicit) objective loss function (see Sect. 2.5) is particularly useful. It can be applied to analyze the sub-optimality of the ranking algorithms which, in practice, may be defined without an explicit probabilistic model. In that sense, our evaluation methodology is more general than reverse engineering, which was used in Szczecinski and Roatis (2022) to unveil the model underlying the FIFA ranking.
Beyond this general recommendation and focusing specifically on improving the FIVB ranking, the following venues can be explored:
  • Considering a time-variant model for the skills in the design of the algorithm, e.g., using ideas already shown before in (Fahrmeir, 1992), (Glickman, 1993), (Knorr-Held, 2000), (Szczecinski and Tihon, 2023). It may require particular attention, as the current version of the FIVB ranking algorithm is not a straightforward implementation of the ML ranking.
  • Analyzing alternative ordinal models McCullagh (1980), taking into account the simplicity of the algorithm they produce, including the use of different CDF in the model (16) (Tutz, 2012, Ch. 9.1.3)(Agresti, 2013, Ch. 8.3), or the Adjacent Categories (AC) models (Tutz, 2012, Ch. 9.4.5) Szczecinski (2022).
@@ Appendix A Notation in FIVB ranking
We show in Table 6, the relationship between our notation and the one used in the FIVB ranking description (FIVB, 2024), where the following abbreviations are used
  • WR: World ranking (here, a FIVB ranking)
@@ Appendix B Optimization of the cross-validation metric
However, the expressions for the gradient and, especially, the Hessian, quickly become cumbersome; see (Burn, 2020). Thus, instead of explicit differentiation, we use the automatic differentiation available in JAX and JAXopt python-compliant packages (Bradbury et al., 2018), (Blondel et al., 2021) with a particularly interesting feature which automatically finds the implicit differentiation required to find the derivative of $\hat{\boldsymbol{\theta}}$ with respect to hyperparameters in $\boldsymbol{p}$ as specified by (72).